OpenStudy (anonymous):

B is the midpoint of A and C. If BC=3x2, and AC=54, find the value of x. Show all work to receive credit. Heres what I got: 3x^2 + 3x^2 = 54 9x + 9x = 54 18x = 54 18x/18 = 54/18 x = 3 3(3)^2 + 3(3)^2 = 54 9^2 + 9^2 = 54 but then it wouldnt equal 54

OpenStudy (yrelhan4):

Eh, 3(3)^2 + 3(3)^2 = 3(9) + 3(9)= 27 + 27 = 54.

OpenStudy (anonymous):

*feels dumb* thanks <3

OpenStudy (yrelhan4):

Ha, ikr. jk.. :P You're welcome.